Week 20: May 11 - 18   Last updated:   14-May-20                                                                          
Maintenance this week Date Time & Length Production resumes                                                                          
520 MeV Friday, May 15 For 6 hr from 08:30 - 14:30, followed by 1 hr start-up, followed by 3 hr Beam Development. 18:30                                                                        
ISAC Shutdown                                                                            
E-LINAC Shutdown                                                                            
ARIEL Shutdown                                                                            
GROUP ACTIVITY   Fault Report Work Permit Completed                                                                    
ATG STF target swap 520 MeV                                                                          
Beamlines SAB3 access to check coring locations for ELSV project 520 MeV                                                                          
Beam Development ISIS studies to commence after I1/ISIS maint. 520 MeV                                                                          
  U-tune studies will occur from 15:30 - 18:30 520 MeV                                                                          
  See the 520 MeV elog for the Beam Dev Schedule.                                                                            
Computing and Networking                                                                              
Controls Investigate ISIS Vacuum Controls issues 520 MeV                                                                          
  Change STF water column height calibration  520 MeV                                                                          
  Add new MUX17 & Test channel  520 MeV                                                                          
  Change B2,18 to MUX 17,0 520 MeV                                                                          
  Install new ccsc.exe (mpmtable) in CCS cluster 520 MeV                                                                          
  Add MUX 17 watchdogs to scan and xtp=-W2 520 MeV                                                                          
Cryogenics and Vacuum Re-tune LINDE (~6 hr) 520 MeV                                                                          
ISIS Regen cryo030 520 MeV                                                                          
  Assist Controls with ISIS vacuum controls troubleshooting 520 MeV                                                                          
  I1 H2 check   520 MeV                                                                          
Operations Sump checks All                                                                          
  SCBCD & SCRM Tests 520 MeV                                                                          
  Search for Cu-LCW (non-active/RF) water leak 520 MeV                                                                          
Remote Handling                                                                              
RF Re-start RF system after LINDE investigations 520 MeV                                                                          
  Water leak check on SAB3 520 MeV                                                                          
Site Services Prep for VFD project electrical work (SAB3 access) [ELSV] 520 MeV                                                                          
Mechanical, Civil/Structural, Electrical, FAS